# HOLE: Homological Observation of Latent Embeddings for Neural Network Interpretability

### v1

@@ 1 Introduction
Deep learning models have gained popularity in recent years [24, 40], and have demonstrated remarkable predictive performance on a wide range of complex tasks, including computer vision [39] and natural language processing [14]. They are often considered to be black boxes, which presents significant challenges, including difficulty in debugging model failures, a lack of trust from end users, potential for perpetuating bias, and regulatory compliance issues [53, 62]. The fundamental problem underlying these issues is the lack of interpretability and explainability—we cannot understand what features the model has learned, how it processes information, or why it makes specific decisions, making it impossible to debug, trust, or verify these systems effectively. This lack of interpretability has profound implications, as there is a rise in the deployment of AI models in critical sectors such as healthcare [59, 69] and finance [34, 45], where transparency, fairness, accountability, and ethical considerations are important [4, 50].
@@ 2 Related Works
Computational topology techniques have gained popularity in data analysis in the last few years, including for understanding neural representations [57].
@@ 2.1 Topological Data Analysis and Visualization
Recent advances in topological data analysis [65] have demonstrated that studying the shape and topological structure of data in high-dimensional spaces provides powerful insights into underlying patterns. Specifically, persistent homology has emerged as a powerful mathematical framework for quantifying the persistence of topological features across multiple scales [17, 76]. The output of persistence homology, called persistence diagrams, is stable under perturbations of the input [12], which opened the door to applications in noisy settings.
The persistence information is visualized using a scatterplot-like visualization, somewhat confusingly called a persistence diagram, or a barchart-like visualization, called a persistence barcode [23], which provides compact representations of the birth and death times of topological features across different scales (1). Vector-space representations such as persistence landscapes [8] have been proposed to transform persistence diagrams into a function-based representation in a vector space, making them usable in machine learning and statistical analysis. Other approaches include persistence images, which create a fixed-size vector representation from diagrams, and specialized kernels [10]. These methods, along with scalable software libraries (e.g., GUDHI [49]), provide practical access to topological methods.
@@
A particularly influential method is Mapper [65], which constructs simplified representations of high-dimensional data by creating nerve complexes of overlapping neighborhoods. Unlike persistent homology, which focuses on quantifying the lifetime of topological features across scales, Mapper addresses the complementary problem of shape summarization—creating interpretable, graph-based visualizations that preserve the topological structure of the data while allowing interactive exploration.
@@ 2.2 Machine Learning Interpretability
The growing complexity of machine learning models has sparked significant interest in interpretability and explainability methods, which can be broadly categorized into local and global approaches. Activation maximization [18] and its variants generate synthetic inputs that reveal preferred stimuli for individual neurons or layers. Local explanation methods focus on understanding individual predictions. LIME (Local Interpretable Model-agnostic Explanations) [60] provides local explanations by learning interpretable models around individual predictions, offering insights into model behavior for specific instances. SHAP (SHapley Additive exPlanations) [48] builds on cooperative game theory to assign importance values to features, providing a unified framework that encompasses several existing explanation methods. Network dissection [5] systematically identifies neurons that correspond to interpretable concepts by measuring alignment with segmentation masks. TCAV (Testing with Concept Activation Vectors) [35] enables testing of user-defined concepts within neural networks, allowing researchers to probe whether models have learned human-interpretable concepts. Global explanation methods aim to understand model behavior more broadly. Feature visualization techniques [56] synthesize inputs that maximally activate specific neurons, revealing what patterns different parts of the network have learned to detect.
@@
Saliency maps represent one of the most widely used approaches for explaining neural network predictions in computer vision. These methods highlight the regions of an input image that most strongly influence the model’s decision. The fundamental approach computes gradients of the output with respect to the input to identify important pixels or regions. Grad-CAM [63] generates visual explanations for CNN decisions by using gradients to produce coarse localization maps highlighting important regions in images. Unlike pixel-level gradient methods, Grad-CAM operates at the feature map level, providing more interpretable visualizations that show which parts of an image the network focuses on for specific predictions. Other gradient-based saliency methods include vanilla gradients, integrated gradients, and guided backpropagation, each offering different perspectives on input attribution. Integrated gradients [67] addressed gradient saturation issues by accumulating gradients along a path from a baseline to the input. SmoothGrad [66] improved saliency map quality by averaging gradients over multiple noisy versions of the input. Layer-wise Relevance Propagation (LRP) [1] proposed a principled approach to decompose predictions into relevance scores for individual features. Occlusion sensitivity analysis [75] provided a complementary approach by systematically blocking image regions to measure their impact on predictions. While saliency maps excel at providing intuitive visual explanations, they primarily focus on input-level attribution and may not capture the complex relationships.
@@
The visualization community has also developed interactive tools for neural network analysis that complement traditional interpretability methods. Summit [31] presents an interactive system for scalable summarization and interpretation of deep neural networks, introducing activation aggregation to discover important neurons and neuron-influence aggregation to identify relationships between neurons. These techniques combine to create attribution graphs that reveal crucial neuron associations across network layers. ActiVis [33] provides visual exploration of industry-scale deep neural network models through coordinated views including computation graph overviews and neuron activation patterns. CNNVis [43] offers interactive visualization of convolutional neural networks, enabling users to explore learned features across different layers through multiple linked views. CNN Explainer [70] provides an educational tool with interactive visualizations that help novices understand the mathematical operations underlying CNNs.
@@ 2.2.1 Topology and Interpretability
During the last few years, there has been a convergence of topological data analysis and deep learning. Hofer et al. [30] showed how persistence can serve as a differentiable layer that preserves topological information in neural networks. Carrière et al. [11] built on this work with PersLay, developing a permutation-invariant architecture that can process persistence diagrams while maintaining their stability properties Moor et al. [54] proposed topology-aware autoencoders that explicitly preserve the topological structure of data in the latent space, which provides explanation as to why the network represents the data in a certain way. These works indicate that several topological approaches can be used to improve the performance of neural networks. Rieck et al. [61] introduced Neural Persistence, which quantifies network complexity by analyzing the lifetime of connections in weight matrices. Improving on the previous work, Watanabe et al. [71] proposed layer-wise topological metrics to understand the complexity of networks at different depths. Purvine et al. [58] conducted detailed empirical studies of convolutional layer activations using topological methods, while Birdal et al. [6] established important connections between persistence and the intrinsic dimension of neural representations. Wheeler et al. [72] developed Activation Landscapes that provide a novel way to visualize and understand training trajectories.
Recent work has also revealed strong connections between topological features and model generalization. Gutierrez et al. [26] demonstrated that persistence statistics can estimate test performance without requiring a validation set, which is a strong indicator that topological features can be used to improve the performance of neural networks. Ballester et al. [2] then showed how these statistics can predict the generalization gap. Collectively, these works strengthen the evidence that topological priors can improve both performance and interpretability—an observation we leverage in the present paper. Rather than instance-level attributions, we pursue a global, layer-by-layer reading of a model’s internal organization—post hoc, comparative, and performed without altering the network. Topology contributes stable, scale-aware cues that make visible where semantic structure emerges, persists, or shifts across layers; we weave these cues into complementary multi-view summaries to support structure-aware explanation.
@@ 3.2 Persistent Homology
Let $X=\{x_{1},x_{2},\ldots,x_{n}\}\subset\mathbb{R}^{d}$ be a finite point cloud (in our case neural network activations). A 0-simplex represents a vertex, a 1-simplex an edge, a 2-simplex a triangle, and so forth. A simplicial complex $K$ is a collection of simplices where every face of a simplex in $K$ is also in $K$ [9]. Formally, a $k$-simplex $\sigma$ is the convex hull of $k+1$ affinely independent points:
@@
where $Z_{k}(K)$ are $k$-cycles (closed chains) and $B_{k}(K)$ are $k$-boundaries. To construct simplicial complexes from point clouds, we utilize the Vietoris-Rips (VR) complex [17]. Given a distance function $d:X\times X\to\mathbb{R}$ and radius $\epsilon\geq 0$, the VR complex is defined as:
@@
Persistent homology is a mathematical framework that allows tracking the evolution of topological features across different scales (i.e., values of $\epsilon$) [9]. The key concept of persistent homology is a filtration, which is a growing sequence of metric balls, used to detect the size and scale of topological features of a dataset [16]. Mathematically, filtration can be defined as a nested sequence of simplicial complexes, $\emptyset=K_{0}\subseteq K_{1}\subseteq K_{2}\subseteq\cdots\subseteq K_{m}$, parameterized by increasing scale parameter $0=\epsilon_{0}<\epsilon_{1}<\epsilon_{2}<\cdots<\epsilon_{m}$. As the filtration parameter $\epsilon$ increases, topological features appear, known as birth, and disappear, known as death. Persistent homology tracks these changes by computing homology groups across the filtration and recording birth-death pairs. Further, a measure, known as persistence, is the difference between the birth and the death values of a feature. Features with high persistence are considered significant topological structures, while short-lived features are typically attributed to noise.
@@ 3.4.1 Task Identification
We identified several tasks and associated visualizations to highlight the features of interest.
[T1] Hierarchy: The hierarchy of ${\mathsf{H}}_{0}$ features helps us better understand the evolution of the class clusters in activation space during the filtration process.
[T2] Separability: Identify the separability of the class clusters in activation space to understand how discriminative they are.
[T3] Homogeneity: While clusters might be well formed, they might also contain multiple classes. We would like to identify non-homogeneous clusters consisting of two or more classes.
[T4] Outliers: The existence and prevalence of outliers may indicate suboptimal model training, suggesting the network has learned non-discriminative representations for some inputs.
@@ 4.1 Implementation
We have implemented HOLE in Python. The hook functions are implemented using PyTorch’s module-level forward hooks, which allow capturing outputs during a forward pass without altering the network architecture. The input data is fed into the neural network, and the activations are recorded at each layer. For a given batch of inputs and a specified network layer, we collect the layer activations and flatten them to form vectors in $\mathbb{R}^{d}$, where $d$ is the number of output units in that layer. The resulting $N\times d$ matrix defines a point cloud of $N$ points in $d$-dimensional space. These probed intermediate activations then act as point clouds for the persistent homology pipeline, which is then used to study the layer topology. We then use GUDHI [49] to compute the persistent homology. Source code is included with our submission for review only, and it will be released as an open-source repository upon publication.
@@ 4.2 Dataset and Network Architectures
The primary dataset used in our experiments is CIFAR-10 [38], a widely adopted benchmark in computer vision for image classification. It consists of 60,000 images at a resolution of 32×32, evenly distributed across 10 mutually exclusive classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. The standard split allocates 50,000 images for training and 10,000 for testing.
We evaluated HOLE for interpretability across multiple neural network architectures to demonstrate its model-agnostic capabilities. ResNet Family: We primarily focus on ResNet-18 as our baseline architecture, with additional experiments on ResNet-34, and ResNet-50 [29] to analyze how network depth affects topological features and interpretability patterns. Vision Transformer (ViT): We ran our experiments on ViT-B/16 and ViT-L/16 models [15], which allowed us to explore transformer architectures.

### v2

@@ 1 Introduction
Deep learning models have gained popularity in recent years [goodfellow2016deep, lecun2015deep], and have demonstrated remarkable predictive performance on a wide range of complex tasks, including computer vision [krizhevsky2012imagenet] and natural language processing [devlin2018bert]. They are often considered to be black boxes, which presents significant challenges, including difficulty in debugging model failures, a lack of trust from end users, potential for perpetuating bias, and regulatory compliance issues [molnar2020interpretable, rudin2019stop]. The fundamental problem underlying these issues is the lack of interpretability and explainability—we cannot understand what features the model has learned, how it processes information, or why it makes specific decisions, making it impossible to debug, trust, or verify these systems effectively. This lack of interpretability has profound implications, as there is a rise in the deployment of AI models in critical sectors such as healthcare [rajkomar2018scalable, topol2019high] and finance [khandani2010consumer, lopez2013machine], where transparency, fairness, accountability, and ethical considerations are important [barocas2016big, mehrabi2021survey].
@@ 2 Related Works
Computational topology techniques have gained popularity in data analysis in the last few years, including for understanding neural representations [purvine2023experimental].
@@ 2.1 Topological Data Analysis and Visualization
Recent advances in topological data analysis [singh2007topological] have demonstrated that studying the shape and topological structure of data in high-dimensional spaces provides powerful insights into underlying patterns. Specifically, persistent homology has emerged as a powerful mathematical framework for quantifying the persistence of topological features across multiple scales [edelsbrunner2002persistence, zomorodian2005computing]. The output of persistence homology, called persistence diagrams, is stable under perturbations of the input [cohen2007stability], which opened the door to applications in noisy settings.
The persistence information is visualized using a scatterplot-like visualization, somewhat confusingly called a persistence diagram, or a barchart-like visualization, called a persistence barcode [ghrist2008barcodes], which provides compact representations of the birth and death times of topological features across different scales (1). Vector-space representations such as persistence landscapes [bubenik2015landscapes] have been proposed to transform persistence diagrams into a function-based representation in a vector space, making them usable in machine learning and statistical analysis. Other approaches include persistence images, which create a fixed-size vector representation from diagrams, and specialized kernels [carriere2017sliced]. These methods, along with scalable software libraries (e.g., GUDHI [maria2014gudhi]), provide practical access to topological methods.
@@
A particularly influential method is Mapper [singh2007topological], which constructs simplified representations of high-dimensional data by creating nerve complexes of overlapping neighborhoods. Unlike persistent homology, which focuses on quantifying the lifetime of topological features across scales, Mapper addresses the complementary problem of shape summarization—creating interpretable, graph-based visualizations that preserve the topological structure of the data while allowing interactive exploration.
@@ 2.2 Machine Learning Interpretability
The growing complexity of machine learning models has sparked significant interest in interpretability and explainability methods, which can be broadly categorized into local and global approaches. Activation maximization [erhan2009visualizing] and its variants generate synthetic inputs that reveal preferred stimuli for individual neurons or layers. Local explanation methods focus on understanding individual predictions. LIME (Local Interpretable Model-agnostic Explanations) [ribeiro2016lime] provides local explanations by learning interpretable models around individual predictions, offering insights into model behavior for specific instances. SHAP (SHapley Additive exPlanations) [lundberg2017shap] builds on cooperative game theory to assign importance values to features, providing a unified framework that encompasses several existing explanation methods. Network dissection [bau2017network] systematically identifies neurons that correspond to interpretable concepts by measuring alignment with segmentation masks. TCAV (Testing with Concept Activation Vectors) [kim2018interpretability] enables testing of user-defined concepts within neural networks, allowing researchers to probe whether models have learned human-interpretable concepts. Global explanation methods aim to understand model behavior more broadly. Feature visualization techniques [olah2017feature] synthesize inputs that maximally activate specific neurons, revealing what patterns different parts of the network have learned to detect.
@@
Saliency maps represent one of the most widely used approaches for explaining neural network predictions in computer vision. These methods highlight the regions of an input image that most strongly influence the model’s decision. The fundamental approach computes gradients of the output with respect to the input to identify important pixels or regions. Grad-CAM [selvaraju2017grad] generates visual explanations for CNN decisions by using gradients to produce coarse localization maps highlighting important regions in images. Unlike pixel-level gradient methods, Grad-CAM operates at the feature map level, providing more interpretable visualizations that show which parts of an image the network focuses on for specific predictions. Other gradient-based saliency methods include vanilla gradients, integrated gradients, and guided backpropagation, each offering different perspectives on input attribution. Integrated gradients [sundararajan2017axiomatic] addressed gradient saturation issues by accumulating gradients along a path from a baseline to the input. SmoothGrad [smilkov2017smoothgrad] improved saliency map quality by averaging gradients over multiple noisy versions of the input. Layer-wise Relevance Propagation (LRP) [bach2015pixel] proposed a principled approach to decompose predictions into relevance scores for individual features. Occlusion sensitivity analysis [zeiler2014visualizing] provided a complementary approach by systematically blocking image regions to measure their impact on predictions. While saliency maps excel at providing intuitive visual explanations, they primarily focus on input-level attribution and may not capture the complex relationships.
Visualization Task Taxonomies
Understanding user analytic activities has been a central concern in information visualization research. Amar et al. [amar2005lowlevel] identified ten low-level analytic primitives (retrieve value, filter, compute derived value, find extremum, sort, determine range, characterize distribution, find anomalies, cluster, and correlate) through empirical analysis of nearly 200 analyst questions across multiple domains. Their taxonomy emphasizes analytic primacy over representational primacy, arguing that visualization systems should be designed to support users’ analytic goals rather than merely presenting data. These primitives serve as foundational building blocks that can be composed into higher-level, domain-specific tasks. Our work builds upon this foundation by composing Amar et al.’s primitives into tasks specifically designed for persistent homology analysis of neural network activations, where tracking multi-scale topological evolution and assessing representation quality require coordinated application of multiple primitive operations.
@@
The visualization community has also developed interactive tools for neural network analysis that complement traditional interpretability methods. Summit [hohman2020summit] presents an interactive system for scalable summarization and interpretation of deep neural networks, introducing activation aggregation to discover important neurons and neuron-influence aggregation to identify relationships between neurons. These techniques combine to create attribution graphs that reveal crucial neuron associations across network layers. ActiVis [kahng2018activis] provides visual exploration of industry-scale deep neural network models through coordinated views including computation graph overviews and neuron activation patterns. CNNVis [liu2017towards] offers interactive visualization of convolutional neural networks, enabling users to explore learned features across different layers through multiple linked views. CNN Explainer [wang2020cnn] provides an educational tool with interactive visualizations that help novices understand the mathematical operations underlying CNNs.
@@ 2.2.1 Topology and Interpretability
During the last few years, there has been a convergence of topological data analysis and deep learning. Hofer et al. [hofer2017deep] showed how persistence can serve as a differentiable layer that preserves topological information in neural networks. Carrière et al. [carriere2020perslay] built on this work with PersLay, developing a permutation-invariant architecture that can process persistence diagrams while maintaining their stability properties Moor et al. [moor2020topoae] proposed topology-aware autoencoders that explicitly preserve the topological structure of data in the latent space, which provides explanation as to why the network represents the data in a certain way. These works indicate that several topological approaches can be used to improve the performance of neural networks. Rieck et al. [rieck2019neural] introduced Neural Persistence, which quantifies network complexity by analyzing the lifetime of connections in weight matrices. Improving on the previous work, Watanabe et al. [watanabe2021topologicalmeasurement] proposed layer-wise topological metrics to understand the complexity of networks at different depths. Purvine et al. [purvine2022cnnactivations] conducted detailed empirical studies of convolutional layer activations using topological methods, while Birdal et al. [birdal2021intrinsic] established important connections between persistence and the intrinsic dimension of neural representations. Wheeler et al. [wheeler2021activationlandscapes] developed Activation Landscapes that provide a novel way to visualize and understand training trajectories.
Recent work has also revealed strong connections between topological features and model generalization. Gutierrez et al. [gutierrez2021phgeneralization] demonstrated that persistence statistics can estimate test performance without requiring a validation set, which is a strong indicator that topological features can be used to improve the performance of neural networks. Ballester et al. [ballester2024generalizationgap] then showed how these statistics can predict the generalization gap. Collectively, these works strengthen the evidence that topological priors can improve both performance and interpretability—an observation we leverage in the present paper. Rather than instance-level attributions, we pursue a global, layer-by-layer reading of a model’s internal organization—post hoc, comparative, and performed without altering the network. Topology contributes stable, scale-aware cues that make visible where semantic structure emerges, persists, or shifts across layers; we weave these cues into complementary multi-view summaries to support structure-aware explanation.
@@ 3.2 Persistent Homology
Let $X=\{x_{1},x_{2},\ldots,x_{n}\}\subset\mathbb{R}^{d}$ be a finite point cloud (in our case neural network activations). A 0-simplex represents a vertex, a 1-simplex an edge, a 2-simplex a triangle, and so forth. A simplicial complex $K$ is a collection of simplices where every face of a simplex in $K$ is also in $K$ [carlsson2009topology]. Formally, a $k$-simplex $\sigma$ is the convex hull of $k+1$ affinely independent points:
@@
where $Z_{k}(K)$ are $k$-cycles (closed chains) and $B_{k}(K)$ are $k$-boundaries. To construct simplicial complexes from point clouds, we utilize the Vietoris-Rips (VR) complex [edelsbrunner2002persistence]. Given a distance function $d:X\times X\to\mathbb{R}$ and radius $\epsilon\geq 0$, the VR complex is defined as:
@@
Persistent homology is a mathematical framework that allows tracking the evolution of topological features across different scales (i.e., values of $\epsilon$) [carlsson2009topology]. The key concept of persistent homology is a filtration, which is a growing sequence of metric balls, used to detect the size and scale of topological features of a dataset [edelsbrunner2010computational]. Mathematically, filtration can be defined as a nested sequence of simplicial complexes, $\emptyset=K_{0}\subseteq K_{1}\subseteq K_{2}\subseteq\cdots\subseteq K_{m}$, parameterized by increasing scale parameter $0=\epsilon_{0}<\epsilon_{1}<\epsilon_{2}<\cdots<\epsilon_{m}$. As the filtration parameter $\epsilon$ increases, topological features appear, known as birth, and disappear, known as death. Persistent homology tracks these changes by computing homology groups across the filtration and recording birth-death pairs. Further, a measure, known as persistence, is the difference between the birth and the death values of a feature. Features with high persistence are considered significant topological structures, while short-lived features are typically attributed to noise.
@@ 3.4.1 Task Identification
We ground our task identification in the low-level analytic task taxonomy established by Amar et al. [amar2005lowlevel], which identified ten primitive operations through empirical analysis of nearly 200 data analysis questions: Retrieve Value, Filter, Compute Derived Value, Find Extremum, Sort, Determine Range, Characterize Distribution, Find Anomalies, Cluster, and Correlate. These primitives form a foundational vocabulary for describing analytic activities in information visualization systems.
Building upon this foundation, we compose four higher-level tasks specifically designed for topological analysis of neural network activations. Each task represents a composition of multiple Amar et al. primitives, applied in the context of persistent homology filtration to address the unique analytical needs of neural network interpretability. Unlike traditional static data analysis, our tasks operate on dynamic topological features that evolve across filtration scales, requiring coordinated application of multiple primitive operations to track hierarchical structure, assess representation quality, and identify anomalous patterns in learned representations.
[T1] Hierarchy: Understand the evolution of class clusters in activation space during the filtration process by tracking how ${\mathsf{H}}_{0}$ components form, persist, merge, and disappear across scales. This task composes Cluster (identify connected components at each filtration level), Characterize Distribution (of component birth/death times), Sort (by persistence values), and Determine Range (of filtration thresholds where meaningful structure emerges). The hierarchical dendrogram reveals multi-resolution organization of learned representations.
[T2] Separability: Assess how discriminatively the network has organized different classes in activation space by measuring the degree of separation between class-specific clusters. This task composes Cluster (identify class groupings), Determine Range (measure inter-cluster distances), Characterize Distribution (of within- vs. between-class distances), and Compute Derived Value (separation metrics at optimal filtration scales). Well-separated clusters indicate effective feature learning.
[T3] Homogeneity: Determine whether topologically coherent clusters correspond to semantically meaningful (single-class) groupings, or whether they incorrectly merge multiple classes. This task composes Cluster (identify persistent components), Retrieve Value (obtain ground-truth class labels for points within each cluster), Compute Derived Value (cluster purity/homogeneity metrics), and Characterize Distribution (of class composition within clusters). Non-homogeneous clusters suggest the network has learned representations that conflate distinct concepts.
[T4] Outliers: Identify data points with anomalous topological behavior—those that remain isolated or merge very late in the filtration—which may indicate poorly learned representations for specific inputs. This task composes Find Anomalies (identify outlying points), Filter (select late-merging or low-persistence components), Retrieve Value (obtain persistence/death times), and Characterize Distribution (of merge times to establish what constitutes “late”). Prevalent outliers suggest suboptimal model training or dataset issues.
@@ 4.1 Implementation
We have implemented HOLE in Python. The hook functions are implemented using PyTorch’s module-level forward hooks, which allow capturing outputs during a forward pass without altering the network architecture. The input data is fed into the neural network, and the activations are recorded at each layer. For a given batch of inputs and a specified network layer, we collect the layer activations and flatten them to form vectors in $\mathbb{R}^{d}$, where $d$ is the number of output units in that layer. The resulting $N\times d$ matrix defines a point cloud of $N$ points in $d$-dimensional space. These probed intermediate activations then act as point clouds for the persistent homology pipeline, which is then used to study the layer topology. We then use GUDHI [maria2014gudhi] to compute the persistent homology. Source code is included with our submission for review only, and it will be released as an open-source repository upon publication.
@@ 4.2 Dataset and Network Architectures
The primary dataset used in our experiments is CIFAR-10 [Krizhevsky2009LearningML], a widely adopted benchmark in computer vision for image classification. It consists of 60,000 images at a resolution of 32×32, evenly distributed across 10 mutually exclusive classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. The standard split allocates 50,000 images for training and 10,000 for testing.
We evaluated HOLE for interpretability across multiple neural network architectures to demonstrate its model-agnostic capabilities. ResNet Family: We primarily focus on ResNet-18 as our baseline architecture, with additional experiments on ResNet-34, and ResNet-50 [he2016deep] to analyze how network depth affects topological features and interpretability patterns. Vision Transformer (ViT): We ran our experiments on ViT-B/16 and ViT-L/16 models [dosovitskiy2020image], which allowed us to explore transformer architectures.